Let P superscript (e) ={p(x)∈R[x]|p(x) = p(−x)} P superscript(o) ={p(x)∈R[x]|p(x) =−p(−x)} a) Check that P superscript (e) and P superscript (o) are subspace of R[x]. b) Show that P superscript (e) =(∑ i a subscript i x superscript i ∈R[x]



a subscript = 0 if i is odd.) P superscript(o) =(∑ i a subscript i x superscript i ∈R[x]



a subscript = 0 if i is even.) Deduce that P superscript (o)∩P superscript (e) ={0}. ( c) Check p(x)+p(−x)∈P superscript(e) for every p(x)∈R(x). Check that the map ψ: R[x]→P superscript(e) given byψ(p(x)) = p(x)+p(−x)/ 2 is a linear map. Further, check that ψ superscript 2 =ψ. Determine the kernel of ψ
1
Expert's answer
2019-03-27T11:39:11-0400

a) If "p_1 (x) \in P^{(e)}" and "p_2 (x) \in P^{(e)}" and if "\alpha, \beta \in \R", then

"p (x) \equiv \alpha p_1 (x) + \beta p_2 (x) = \alpha p_1 (- x) + \beta p_2 (- x) = p (-x) ,"

so that "p (x) \in P^{(e)}". Similarly, if "p_1 (x) \in P^{(o)}" and "p_2 (x) \in P^{(o)}", then

"p (x) \equiv \alpha p_1 (x) + \beta p_2 (x) \\ {} = - \alpha p_1 (- x) - \beta p_2 (- x) = - p (-x) \, ,"

so that "p (x) \in P^{(o)}". This means that "P^{(e)}" and "P^{(o)}" are linear subspaces of "\R [x]".


b) Let "p (x) = \sum_i a_i x^i \in P^{(e)}". Then "p(x) = p(-x)", so that

"\sum_i a_i x^i = p(x) = \frac12 \left[ p(x) + p (-x) \right] \\ {} = \frac12 \left[ \sum_i a_i x^i + \sum_i a_i (- 1)^i x^i \right] \\ {} = \frac12 \sum_i a_i \left[ 1 + (-1)^i \right] x^i \, ."

Equating the coefficients on the left-hand side and on the right-hand side, we have

"a_i = \frac12 a_i \left[ 1 + (-1)^i \right] \, ,"

placing no restriction on "a_i" if "i" is even, and implying "a_i = 0" if "i" is odd. Hence,

"P^{(e)} = \left\{ \sum_i a_i x^i \in \R [x] \, , a_i = 0 ~ \text{if}~ i ~ \text{is odd} \right\} \, ."


Let "p (x) = \sum_i a_i x^i \in P^{(o)}". Then "p (x) = - p (-x)", so that

"\sum_i a_i x^i = p(x) = \frac12 \left[ p(x) - p (-x) \right] \\ {} = \frac12 \left[ \sum_i a_i x^i - \sum_i a_i (- 1)^i x^i \right] \\ {} = \frac12 \sum_i a_i \left[ 1 - (-1)^i \right] x^i \, ."

Equating the coefficients on the left-hand side and on the right-hand side, we have

"a_i = \frac12 a_i \left[ 1 - (-1)^i \right] \, ,"

placing no restriction on "a_i" if "i" is odd, and implying "a_i = 0" if "i" is even. Hence,

"P^{(o)} = \left\{ \sum_i a_i x^i \in \R [x] \, , a_i = 0 ~ \text{if}~ i ~ \text{is even} \right\} \, ."

The space "P^{(e)} \cap P^{(o)}" consists of "p (x) = \sum_i a_i x^i \in \R[x]" such that "a_i = 0" for "i" both odd and even, hence, for all "i". Thus, "P^{(e)} \cap P^{(o)} = \{ 0 \}".


c) We have "q (x) = p (x) + p (-x) = p (-x) + p(x) = q (-x) \in P^{(e)}" for every "p (x) \in \R [x]".


If "p_1 (x), p_2 (x) \in \R [x]" and "\alpha , \beta \in \R", then

"\psi \left( \alpha p_1 (x) + \beta p_2 (x) \right) \\ {} = \frac12 \left( \alpha p_1 (x) + \beta p_2 (x) + \alpha p_1 (-x) + \beta p_2 (-x) \right) \\ {} = \frac12 \left( \alpha p_1 (x) + \alpha p_1 (-x) \right) + \frac12 \left( \beta p_2 (x) + \beta p_2 (-x) \right) \\ {} = \alpha \frac12 \left( p_1 (x) + p_1 (-x) \right) + \beta \frac12 \left( p_2 (x) + p_2 (-x) \right) \\ {} = \alpha \psi \left( p_1 (x) \right) + \beta \psi \left( p_2 (x) \right) \, ."

This demonstrates that "\psi \left( p (x) \right)" is a linear map.


Further, we have

"\psi^2 \left( p(x) \right) = \psi \left( \frac12 \left( p(x) + p(-x) \right) \right) \\ {} = \frac12 \left( \frac12 \left( p(x) + p(-x) \right) + \frac12 \left( p(-x) + p(x) \right) \right) \\ {} = \frac12 \left( p(x) + p(-x) \right) = \psi \left( p(x) \right) \, ."

Thus, "\psi^2 = \psi".


The kernel of "\psi" is the subspace formed by "p (x) \in \R[x]" such that "\psi \left( p (x) \right) = 0." The last condition is equivalent to "p (x) + p (-x) = 0", or "p (-x) = - p (x)". Hence, the kernel of "\psi" is the subspace "P^{(o)}", "{\rm ker}\, \psi = P^{(o)}".